OpenStudy (anonymous):

Can someone help me with this please S = n/2 (a+t) solve for n Thanks

OpenStudy (retireed):

Yes

OpenStudy (retireed):

Is this the equation ? \[S = (n/2) (a+t)\]

OpenStudy (retireed):

I'll assume so... divide both sides by (a+t) S / (a+t) = n/2 multiply both sides by 2 2S / (a + t) = n n = 2S / (a + t)

OpenStudy (anonymous):

Thanks it was